HOT-119: Chief Scientist Report


Chief Scientist: D. HEBEL



HOT 119 Cruise Report
R/V Kaimikai O Kanaloa
16-20 Oct., 2000
HALE ALOHA Recovery
(HA8B)
14-15 Oct., 2000



Personnel List
HALE ALOHA Recovery

JGOFS group:
Dale Hebel                      Scientist (co-PI  JGOFS)                UH 
Lance Fujieki                   Computer Specialist                     UH
Terrance Houlihan               Volunteer (Chief Scientist)             MOBY 
Chuck Stump                     Scientist                               UW
Matthew Erickson                Research Associate                      UH

WOCE group:
Mark Valenciano                 Electronic Technician                   UH

HOT 119:
 
JGOFS group:
Dale Hebel                      Chief Scientist (co-PI JGOFS)           UH
Lance Fujieki                   Computer Specialist                     UH
Colleen Allen                   Research Associate                      UH
Matt Church                     Visiting Graduate Student               UH
Ann Gasc                        Scientist                               UH
Matthew Erickson                Research Associate                      UH
Karin Bjorkman                  Scientist                               UH
Ursula Magaard                  Research Associate                      UH

WOCE group:
Fernando Santiago-Mandujano*    Research Associate                      UH
Lal Ratnapala                   Graduate Assistant                      UH
Mark Valenciano                 Electronic Technician                   UH
Jeremiah Johnson                Research Associate                      UH
Javier Mendez-Nuarez            Volunteer                               UH

Associated projects:
Tom Gregory                     Research Associate                      UH
Chuck Stump                     Scientist                               UW

STAG:
Steve Poulos                    Electronic Technician                   UH-UMC
Dave Gravatt                    Deck Technician                         UH-UMC

*Watch Leader




Events log (approximate HST):
Saturday, 14 Oct.
2300    Departed Snug Harbor

Sunday, 15 Oct.
0046    c/c 296 degrees for Kauai
0715    Sighted buoy
0750    All stop at buoy
0756    Transducer in the water
0808    Transducer on deck, no signal
0836    Avon in the water with R. Barnes and C. Gutzeit
0900    Both anemometers and antenna removed from buoy
0905    Tow line on buoy
0916    Tag lines on buoy
0922    Buoy on deck
0929    Avon on deck
1040    Recovery completed
1048    Underway course 117 T
2050    Arrived Snug Harbor, began offloading buoy and equipment 


Monday, 16 Oct.
0900    Departed Snug Harbor
0930    Fire/abandon ship drill, science meeting    
1150    Arrived Kahe Pt. (Sta. 1)
1205    Weight cast (1000 m)
1250    PRR/TSRB casts 
1340    s1c1
1440    Departed Kahe

Tuesday, 17 Oct.
0000    Arrived Station ALOHA (Sta. 2)
0025    s2c1
0035    CTD on deck
0040    Net tow
0105    Net tow
0145    Began sediment trap deployment
0230    Completed sediment trap deployment (22 44.47'N, 157 59.17'W)
0350    s2c2 (WOCE deep)
0535    s2c2 retrieving from 4770 db
0720    s2c2 on deck
0910    s2c3 (WOCE shallow) 
1010    Net tow
1040    Net tow
1120    PRR/TSRB cast
1200    s2c4
1305    Net tow
1335    Net tow
1400    in situ pumping  
1520    s2c5
1805    s2c6 
2105    s2c7
2210    Net tow 
2250    Net tow

Wednesday, 18 Oct.
0005    s2c8
0100    Net tow
0150    Net tow
0205    G.O. cast (25&45m external closing)     
0300    s2c9
0600    Deployed primary productivity array (22 45.4'N, 157 58.9'W)
0610    s2c10
0905    s2c11
1000    Net tow
1030    Net tow
1120    PRR-600/TSRB cast
1200    s2c12
1300    Net tow
1400    in situ pump
1500    s2c13
1800    Recovered PP array
1810    s2c14 
2100    s2c15   

Thursday, 19 Oct.
0000    s2c16 (WOCE deep-2)
0345    Transit sediment traps
0720    Recovered sediment traps (location ??)
1010    Arrive HALE ALOHA
1035    Transducer deployed and signal received from releases
1100    s8c1
1240    Transit Kaena Pt.
1750    s6c1
1840    CTD at 2490 m
1930    CTD on deck and began Honolulu transit

Friday 20 Oct.
0700    Arrived Snug Harbor
 


HALE ALOHA Recovery Narrative:

HOT 119 deja vu.  It has not been more than 8 mos earlier that we were
scheduled to go out on HOT 111.  While loading we were informed by, D.
Karl, that he suspected that our deep-sea mooring had broken free and
was adrift.  At that time we assessed the available information and
determined that the buoy was indeed adrift and immediately planned a
recovery operation.  Unfortunately, the ship, which had just received
an upgrade of the SCR engine drive system was experiencing
difficulties.  These difficulties translated into a stream of delays
which encompassed the better part of a week with no immediate
resolution.  During this time we monitored the drift of our way-ward
buoy via Argos positioning and devised a recovery plan that utilized
the UNOLS ship R/V Thomas Thompson which had just completed operations
in Hawaiian waters.  In this regard HOT 119 began with a similar
scenario.  Again we were mobilizing for a HOT cruise when we discovered
that our buoy had broke free, and in a similar situation, the ship
personnel had discovered a problem with the ship which would delay our
departure (this time a hole in the hull).  The repairs were forecasted
at approx. 1 week, therefore we looked into an alternative recovery
vessel.  Due to the forecasted high sea-state these plans were
rejected.  Fortunately, the  repairs proceeded ahead of schedule and it
was determined that we could launch a recovery operation on the KOK
before the official HOT cruise which had been rescheduled for 16-20
Oct.  The recovery was scheduled for the 14-15 Oct. the weekend prior
to the scheduled HOT cruise.

The recovery operation (HA8B) departed on Saturday night to arrive off
the southeaster side of Kaui in the early morning to intercept the
drifting buoy.  Although the seas were forecast in the 18' range, upon
arrival we found 4-6' seas although the winds were steadily increasing
from less than 20 kts to upwards of 30 kts.  During this period the
captain launched a small boat operation to remove valuable
instrumentation from the superstructure of the buoy and secure the buoy
for retrieval.  This action, in addition to the skilled abilities of
STAG deck technician Dave Gravatt resulted in a near perfect
retrieval.  Additional deck support was provided by Terrance Houlihan ,
whom had recently transferred to another project and Chuck Stump from
the Univ. of Washington Seattle.  Through these efforts we were able to
recover all instrumentation with the exception of 2 Seatcats below the
point where the cable separated.  Unfortunately, the MBARI nitrate
analyzers and OSU optical buoy where badly damaged.

Upon retrieval it was apparent that the mooring line had severed at
approximately 560 m where an instrument was mounted on the 5/16 plastic
jacketed steel cable.  At the time of retrieval it appeared that in the
area where the cable parted the upper Seacat and thermistor (?) slid
down the cable (more likely pulled down) to the location of the lower
Seacat.  At the location of the lower Seacat the cable was tightly
twisted around the mounting bracket and it was here that the cable
parted.  Later at UH microscopic examination of the area where the
cable parted suggested that the parting was not the result of an
episodic event (although we did find leader material and components
which appeared to be from a long-line array on the retrieved mooring
line).  Rather, abrasion of the cable.  At this time, I understand that
the severed section of the cable has been sent out for professional
examination.  Following the recovery we returned to Snug Harbor and
offloaded equipment that night in preparation for the followi

HOT 119 Narrative:

The following morning HOT 119 departed aboard the R/V Kaimikai O
Kanaloa (KOK).  Captain Hayes was the master of the vessel and Dale
Hebel chief scientist.  There was a total of 17 participants in the
scientific party composed of 5 WOCE, 8 JGOFS, 2 ancillary investigators
and 2 STAG. We departed Snug on 16 Oct., occupying stations at Kahe Pt.
(sta. 1), Station ALOHA (sta. 2), HALE ALOHA (sta. 8) and Kaena Pt.
(sta. 6).  All scheduled work was completed and all samples collected.
CTD operations were conducted at stations 1, 2, 8 and 6.  One ~1000 m
CTD cast was conducted at stations 1&8; 1 <250m, 14 ~1000 m, and 2
~4800 m CTD casts at Station ALOHA; one ~2500m CTD cast at Kaena Pt.
(sta.6).  Other over-the-side operations included 3 light casts, 13 net
tows, 2 in situ pumping operations, 1 G.O. cast, floating sediment
traps and primary productivity array.  All operations were routine with
the exception of an external closing General Oceanics bottle primary
productivity experiment and addition of sta. 6. The underway/continuous
thermosalinograph, ADCP, and fluorometer were operable and functioned
properly. WOCE met. Obs and limited ship met. data were collected as
well as aerosol measurements.  Overall the weather was partly cloudy
with

At Sta. ALOHA the day-day cruise schedule was similar to a generic HOT
cruise (eg., see HOT 111 cruise report). However on HOT 119 we occupied
sta. 8 (HALE ALOHA) even though our mooring was absent marking a change
in protocol to include sta. 8 as a regular station to assess short
spacial variability and again occupied sta. 6 (Kaena Pt.) on the return
leg.





Weather
HALE ALOHA Recovery (Oct. 15) & HOT 119 (Oct. 16-20):

The weather was mostly cloudy skies with moderate seas and wind.  Below
is listed the cruise bridge log descriptions and the various values
representing the range for that day.  Under wind, sea, and swell there
will be two designations, the first is the direction (in degrees), the
second for wind is in kts, sea in Beauford force, and swell in feet,
barometer in inches of Hg, temp F (dry bulb) and clouds in tenths.

Day/Date       Wind           Sea          Swell        Barometer    Temp   Clouds

Sun 15 Oct.    030-090,10-25  030-090,2-4  000-330,3-4  29.92-30.02  77-80  5-9
Mon 16 Oct.    070-090,4-24   070-090,2-4  070-110,2-8  29.90-29.98  77-88  5-8
Tues 17 Oct.   090-100,17-19  090-100,3    090,3-4      29.94-30.01  75-83  4-8
Wed 18 Oct.    060-100,15-20  060-100,3    090/330,4    29.94-30.04  74-82  6-9
Thur 19 Oct.   070-090,13-24  070-090,3-4  090,4-6      29.95-30.04  76-81  3-9
Fri. 20 Oct.*  080-100,20-22  080-100,3-4  120-130,3-4  29.98-30.00  72-73  2-5

*Two entries  (0200 & 0600 hrs)


Equipment and methods:  All standard equipment functioned properly and
all methodology was standard.  This was the first cruise to use the new
CTD winch wire.  The wire was lubricated during the second WOCE deep
cast.  Also, the acoustic releases were successfully interrogated at
the HALE ALOHA site.
                 

Sub component programs:

Investigator:                   Project:
-----------------               ----------
Bob Bidigare                    HPLC pigments/UH
Michael Landry                  zooplankton dynamics/UH

Ancillary programs:

Investigator:                   Project:
-----------------               ----------
Charles Keeling                 CO2 dynamics and intercalibration/SIO
Paul Quay                       DIC and 13C/UW
Steve Emerson                   O2, N2, Ar, dynamics
John Porter                     aerosols/UH
Abbott/Letelier                 optical measurements/OSU
Claudia Benitz-Nelson           phosphorus isotopes, Th234/UH

Students:
Matt Church                     DOC, Archea dynamics/UH

Others:
Hebel, Dore, Karl               EOC, 1 prod. comparison/UH
Karin Bjorkman                  phosphorus experiments/UH
John Dore                       phycoerythrin distributions, 
                                nitrification rates/UH


Notable events:
1.      Addition of sta. 6
2.      First cruise to use new CTD wire 
3.      Lubrication of new CTD wire
4.      Interrogation of HA acoustic releases
5.      Recovery of drifting HALE ALOHA buoy and instrumentation